OpenStudy (anonymous):

Calculas: Any and all help is appreciated!!!!!? Consider the curve given by x2 + sin(xy) + 3y2 = C, where C is a constant. The point (1, 1) lies on this curve. Use the tangent line approximation to approximate the y-coordinate when x = 1.01. 0.996 1 1.004 Cannot be determined 1.388

OpenStudy (shamil98):

calculas? i've never heard of it

OpenStudy (shamil98):

i joke.

OpenStudy (anonymous):

can u help?????

OpenStudy (shamil98):

the tangent line is the derivative find the first derivative of x2 + sin(xy) + 3y2 = C

OpenStudy (anonymous):

1.004?????

OpenStudy (anonymous):

hold on

OpenStudy (shamil98):

find the first derivative and then solve for y when x = 1.01, that's what they're asking right?...

OpenStudy (anonymous):

d/dx(x^2+sin(xy) + 3y2=((2)(x)^(2-1))+sin'(xy)+(2)(3)y^(2-1))+????? is that right?

OpenStudy (shamil98):

@Isaiah.Feynman feel free to butt in , i suck at calc >.>

OpenStudy (anonymous):

y cos(xy)+2x

OpenStudy (shamil98):

2x + cos (xy) +6yy' = 0 i think you're supposed to use the product rule for sin (xy) tho...

OpenStudy (anonymous):

OK

OpenStudy (anonymous):

So...

OpenStudy (shamil98):

\[\huge \frac{ d }{ dx } \sin(xy)\] ..

OpenStudy (anonymous):

(2x)+(cos(xy))+(6y)

OpenStudy (anonymous):

2(1.01)+cos(1.01)+6y=0

OpenStudy (anonymous):

2(1.01)+cos(1.01)y+6y=0

OpenStudy (anonymous):

2(1.01)+0.531861y+6y=0

OpenStudy (anonymous):

6.531861y=2.02

OpenStudy (anonymous):

y=(6.531861/2.02)

OpenStudy (anonymous):

y=3.2335945544554455445544554455445544554455445544554455

OpenStudy (anonymous):

idk